# MetaCluster: Enabling Deep Compression of Kolmogorov-Arnold Network

### v1

@@ Abstract
Kolmogorov-Arnold Networks (KANs) replace scalar weights with per-edge vectors of basis coefficients, thereby boosting expressivity and accuracy but at the same time resulting in a multiplicative increase in parameters and memory. We propose MetaCluster, a framework that makes KANs highly compressible without sacrificing accuracy. Specifically, a lightweight meta‑learner, trained jointly with the KAN, is used to map low‑dimensional embedding to coefficient vectors, shaping them to lie on a low‑dimensional manifold that is amenable to clustering. We then run K-means in coefficient space and replace per‑edge vectors with shared centroids. Afterwards, the meta‑learner can be discarded, and a brief fine‑tuning of the centroid codebook recovers any residual accuracy loss. The resulting model stores only a small codebook and per-edge indices, exploiting the vector nature of KAN parameters to amortize storage across multiple coefficients. On MNIST, CIFAR-10, and CIFAR-100, across standard KANs and ConvKANs using multiple basis functions, MetaCluster achieves a reduction of up to $80\times$ in parameter storage, with no loss in accuracy. Code will be released upon publication.
@@ 1 Introduction
Kolmogorov–Arnold Networks (KANs) have recently emerged as a compelling alternative to multi‑layer perceptrons (MLPs), delivering strong results in equation modeling and scientific machine learning, often with improved task performance at comparable or lower parameter counts in those settings (Liu et al., 2024; Li et al., 2025; Coffman & Chen, 2025; Koenig et al., 2024). Recently, KANs have also begun to show promise in computer vision (Yang & Wang, 2024; Raffel & Chen, 2025). However, unlike equation modeling, at larger scales, KANs frequently incur a substantial parameter overhead relative to MLPs (Yu et al., 2024). This overhead stems from KANs per-edge degrees of freedom: each connection carries a vector of basis coefficients (e.g., B-spline weights) rather than a single scalar weight, resulting in a multiplicative increase in parameters.
@@
One approach that targets this multiplicative increase in parameters is weight sharing, which clusters parameters into a codebook that stores compact indices. Unfortunately, naively applying weight sharing to KANs is ineffective. Instead of clustering scalars (as in MLPs), we must cluster high‑dimensional coefficient vectors in KANs. In such high-dimensional spaces, absolute distances grow but also concentrate (nearest and farthest become similar). With such an effect brought by the curse of dimensionality, typical clustering methods struggle to form tight clusters (Beyer et al., 1999; Donoho et al., 2000).
@@
We address this challenge with MetaCluster, a three‑stage compression framework that merges meta‑learning with weight sharing. First, a small meta-learner maps low-dimensional embeddings into per-edge coefficient vectors, constraining KAN activations to a low-dimensional manifold while training on the task loss. This manifold shaping makes the coefficient vectors highly clusterable. Second, we run K-means on the generated coefficients, replacing per‑edge weights with codebook centroids indexed with compact codes. Finally, we discard the meta‑learner and embeddings, and lightly fine‑tune the centroids to recover any accuracy loss. Since each centroid stores an entire coefficient vector, the codebook amortizes over many scalars, yielding a much higher compression factor for KANs than for MLPs at the same number of clusters.
@@
We validate MetaCluster on two model families consisting of a fully-connected KAN (Liu et al., 2024) and a convolutional KAN (ConvKAN) (Bodner et al., 2024; Drokin, 2024). For each of these models, we test the efficacy of using B-Splines, radial basis functions (RBFs), and Gram polynomials as the bases (Liu et al., 2024; Li, 2024; SS et al., 2024). To further verify the robustness of our approach, we validate it across MNIST (Deng, 2012), CIFAR-10, and CIFAR-100 (Krizhevsky, 2009). From our experiments, we find that MetaCluster achieves a reduction of up to $80\times$ in parameter storage relative to the uncompressed KAN without degrading accuracy, is robust across various architectures and datasets, and ablations confirm that enforcing a low-dimensional manifold is key to high-quality clustering.
@@
    We identify the potential of weight-sharing for reducing the KANs memory footprint and, to our knowledge, provide the first effective weight‑sharing method tailored to KANs.
@@
    We provide extensive experiments demonstrating up to $80\times$ memory reduction with no loss in accuracy, along with extensive ablations.
@@ 2 Preliminaries and Motivation
2.1 Kolmogorov-Arnold Network
KANs have gained attention as an alternative to conventional MLPs (Liu et al., 2024). Their design is motivated by the Kolmogorov–Arnold representation theorem, which guarantees that any continuous multivariate function on a bounded domain, $f:[0,1]^{n}\rightarrow\mathbb{R}$, can be expressed as a finite sum of compositions of univariate continuous functions, $\phi_{q,p}:[0,1]\rightarrow\mathbb{R}$ and $\phi_{q}:\mathbb{R}\rightarrow\mathbb{R}$ such that
@@ 2.1 Kolmogorov-Arnold Network
While Equation 1 captures the classical theoretical form, Liu et al. (2024) demonstrates a practical generalization that permits arbitrary width and depth tailored to the task. We can express such a formulation with $L$ layers as
$f(\mathbf{x})=(\Phi_{L}\circ\Phi_{L-1}\circ...\circ\Phi_{1})\mathbf{x},\qquad\Phi_{l}=\begin{bmatrix}\phi_{l,1,1}(\,\cdot\,)&\cdots&\phi_{l,1,n_{l}}(\,\cdot\,)\\ \vdots&\ddots&\vdots\\ \phi_{l,n_{l+1},1}(\,\cdot\,)&\cdots&\phi_{l,n_{l+1},n_{l}}(\,\cdot\,)\end{bmatrix}$ (2)
In Equation 2, we let $n_{l}$ denote the number of inputs to layer $l$, with inputs $x_{l,i}$. Each edge from input $i$ to output $j$ in layer $l$ is equipped with a learnable univariate activation $\phi_{l,j,i}(\cdot)$, for $i\in[1,n_{l}]$ and $j\in[1,n_{l+1}]$.
@@
From its increased representation power compared to the MLP, the KAN has offered impressive results on equation modeling tasks for scientific applications (Li, 2024; Li et al., 2025; Coffman & Chen, 2025; Koenig et al., 2024) and has even started to extend its reach into computer vision (Raffel & Chen, 2025; Yang & Wang, 2024). Despite these gains, widespread adoption in modern large‑scale architectures has been hindered by memory inefficiency (Yu et al., 2024). The root cause is structural: each KAN edge carries a vector of basis coefficients (e.g., B‑spline weights), whereas an MLP edge carries a single scalar. Ignoring biases, an arbitrary MLP with $L$ layers will possess $\sum_{l=0}^{L-1}(n_{l}\times n_{l+1})$ parameters, whereas a KAN will possess $\sum_{l=0}^{L-1}(n_{l}\times n_{l+1})\times(|\mathbf{w}|)$ parameters. Thus, for identical topologies, a KAN is approximately $|\mathbf{w}|$ times larger than its MLP counterpart.
2.2 Motivations and Challenges of Weight Sharing for KAN
These observations motivate a compression strategy that targets the dimensionality of the vector of basis coefficients. Weight sharing is one method that reduces dimensionality directly by clustering parameters into a small codebook and storing compact indices. In its classical form for MLPs, one applies K‑means to the set of scalar weights $\mathbf{W}={w_{1},w_{2},\dots,w_{n}}$, obtaining centroids $\mathbf{C}={c_{1},c_{2},\dots,c_{k}}$ and assignments that minimize within‑cluster sum of squares (Han et al., 2015):
@@ 2.2 Motivations and Challenges of Weight Sharing for KAN
While weight sharing is promising to save substantial storage for KANs, directly applying it to KANs for model compression is nontrivial. For instance, since each edge on a KAN is defined by $|\mathbf{w}|$ weights rather than $1$ as in the MLP, the dimensionality of the points we must cluster is far greater. In increasing the dimensionality, the vector space becomes sparser and points become more equally spaced apart, making clustering points more difficult (Beyer et al., 1999; Donoho et al., 2000). Therefore, to effectively cluster the KAN activations, we require a fundamentally different method for transforming the high-dimensional vector into a more manageable space.
@@ 3 Methods
This section presents MetaCluster, a three‑stage framework that makes per‑edge KAN activation coefficients amenable to clustering and compression. First, we train a single meta-learner that turns compact embeddings into full activation‐coefficient vectors, forcing them onto a task-aligned low-dimensional manifold. Next, we run K-means on those vectors and record, for each edge, which centroid it belongs to in that layer. Finally, we discard the meta-learner and embeddings, replace per-edge weights with their assigned centroids via a lookup table, and briefly fine-tune the network to recover any lost accuracy.
@@ 3.1 Manifold Learning through a Meta-learner
To demonstrate the manifold‑shaping effect of the meta‑learner, we train a two‑layer KAN with a B‑spline basis on CIFAR-10 Krizhevsky (2009) for one epoch under three settings: (i) a meta‑learner with $d_{\text{emb}}=1$, (ii) a meta‑learner with $d_{\text{emb}}=2$, and (iii) a baseline KAN without a meta‑learner. We collect all per‑edge activation coefficient vectors $\mathbf{w}_{i}$} and visualize them with t‑SNE. Each point corresponds to $\mathbf{w}_{i}\in\mathbb{R}^{|\mathbf{w}|}$, where $|\mathbf{w}|=G+k+1$ with $G=5$ and $k=3$, yielding $|\mathbf{w}|=9$. As shown in Figure 1(a), when $d_{\text{emb}}=1$ the model organizes coefficients along an effectively one‑dimensional manifold. Increasing to $d_{\text{emb}}=2$ produces the well‑structured two‑dimensional sheet shown in Figure 1(b), indicating that the meta‑learner concentrates variability onto task‑relevant directions. In contrast, the baseline KAN provided in Figure 1(c) yields a diffuse cloud without coherent low‑dimensional structure, which hampers downstream clustering. These visualizations substantiate our premise: shaping coefficient vectors via a low‑dimensional embedding makes them markedly more clusterable, a property we later translate into the higher compression at fixed accuracy shown in Section 4.3.
@@
3.2 KAN Activation Compression with K-means MetaClustering
@@ 3.4 Storage Analysis
As in Equation 5, $n$ denotes the number of connections, $k$ the number of centroids, and $b$ the number of bits used to represent each edge. Notably, the term $|\mathbf{w}|$ in the denominator reduces the relative contribution of $n\log_{2}(k)$ to the overall storage. Intuitively, this happens because each centroid stores more weight information ($|\mathbf{w}|$ entries per centroid), so the overhead of indexing and storing the centroids becomes proportionally smaller. As a result, KAN benefits more from the compression scheme than a standard MLP.
@@ 4 Experiments
To evaluate the capabilities of MetaCluster, we conduct extensive experiments and an ablation study on both fully-connected and convolutional architectures. In Section 4.1, we report the experimental setup, and in Section 4.2, we report the results of the fully-connected and convolutional architecture on MNIST, CIFAR10, and CIFAR100 (Deng, 2012; Krizhevsky, 2009). Then, in Section 4.3, we provide the ablation study for our approach, covering the influence of the KAN meta-learner embedding size, basis coefficient vector sizes, and the cluster count on CIFAR10 (Krizhevsky, 2009).
4.1 Experimental Setup
For all our experiments, we did a 90/10 training/validation split of the training data. We apply data augmentation to the train set for the convolutional architecture, following an identical procedure to Drokin (2024). We compare our approach across 24 model schemes. The base model for each scheme consists of KAN (B-spline basis) (Liu et al., 2024), FastKAN (RBF basis) (Li, 2024), and GramKAN (Gram polynomial basis)(Drokin, 2024). In our naming conventions, we include Cluster to designate a clustered model, Meta to designate a model using a meta-learner, a design that originates from Zhao et al. (2025), and Conv to designate a model with a convolutional architecture (Bodner et al., 2024; Drokin, 2024).
4.1.1 Fully-Connected Setup
Our fully-connected architecture follows Zhao et al. (2025) in that we stack two fully-connected KAN layers, with a 32‐dimensional hidden state between them. Each KAN variant uses a SiLU activation (Elfwing et al., 2018). The meta-learner variants contained a meta-learner with a hidden dimension of 32. Concretely:
  • KAN: degree–3 B‐splines, grid range [1,1][-1,1], grid size 55 (Liu et al., 2024).
  • FastKAN: 88 RBFs over [2,2][-2,2] (Li, 2024).
  • GramKAN: degree–3 polynomial (Drokin, 2024).
We train with AdamW (Loshchilov & Hutter, 2017) and for at most 50 epochs, with early stopping (patience = 10). After the final epoch, we cluster the learned weights into 16 groups, then fine‐tune for an additional 5 epochs.The full list of hyperparameters is given in Appendix A.1. We report performance in terms of classification accuracy and model memory footprint.
4.1.2 Convolutional Setup
  • KANConv: degree–3 B‐splines, grid range [3,3][-3,3], grid size 55 (Liu et al., 2024).
  • FastKANConv: 88 RBFs on [3,3][-3,3] (Li, 2024).
  • GramKANConv: degree–3 polynomial (Drokin, 2024).
Training again uses AdamW (Loshchilov & Hutter, 2017) for up to 150 epochs with early stopping (patience = 10). Final weights are clustered into 16 groups and fine‐tuned for 5 epochs. See Appendix A.1 for the complete hyperparameter sweep. Evaluation metrics are classification accuracy and model memory footprint.
4.2 Results

### v2

@@ Abstract
Kolmogorov-Arnold Networks (KANs) replace scalar weights with per-edge vectors of basis coefficients, thereby increasing expressivity and accuracy while also resulting in a multiplicative increase in parameters and memory. We propose MetaCluster, a framework that makes KANs highly compressible without sacrificing accuracy. Specifically, a lightweight meta-learner, trained jointly with the KAN, maps low-dimensional embeddings to coefficient vectors, thereby shaping them to lie on a low-dimensional manifold that is amenable to clustering. We then run K-means in coefficient space and replace per-edge vectors with shared centroids. Afterwards, the meta-learner can be discarded, and a brief fine-tuning of the centroid codebook recovers any residual accuracy loss. The resulting model stores only a small codebook and per-edge indices, exploiting the vector nature of KAN parameters to amortize storage across multiple coefficients. On MNIST, CIFAR-10, and CIFAR-100, across standard KANs and ConvKANs using multiple basis functions, MetaCluster achieves a reduction of up to $80\times$ in parameter storage, with no loss in accuracy. Similarly, on high-dimensional equation modeling tasks, MetaCluster achieves a parameter reduction of $124.1\times$, without impacting performance. Code will be released upon publication.
@@ 1 Introduction
Kolmogorov–Arnold Networks (KANs) have recently emerged as a compelling alternative to multi-layer perceptrons (MLPs), delivering strong results in equation modeling and scientific machine learning, often with improved task performance at comparable or lower parameter counts in those settings (Liu et al., 2024b; Li et al., 2025; Coffman and Chen, 2025; Koenig et al., 2024). Recently, KANs have also shown promise in computer vision (Yang and Wang, 2024; Raffel and Chen, 2025). However, unlike equation modeling, KANs at larger scales often incur substantial parameter overhead relative to MLPs (Yu et al., 2024). This overhead stems from the per-edge degrees of freedom in KANs: each connection carries a vector of basis coefficients (e.g., B-spline weights) rather than a single scalar weight, resulting in a multiplicative increase in the number of parameters.
@@
One approach that targets this multiplicative increase in parameters is weight sharing, which clusters parameters into a codebook that stores compact indices. Unfortunately, naively applying weight sharing to KANs is ineffective. Instead of clustering scalars (as in MLPs), we must cluster high-dimensional coefficient vectors in KANs. In such high-dimensional spaces, absolute distances increase while also concentrating (the nearest and farthest become similar). Given the curse of dimensionality, typical clustering methods struggle to form tight clusters (Beyer et al., 1999; Donoho and others, 2000).
@@
We address this challenge with MetaCluster, a three-stage compression framework that merges meta-learning with weight sharing. First, a small meta-learner maps low-dimensional embeddings to per-edge coefficient vectors, thereby constraining KAN activations to a low-dimensional manifold during training on the task loss. This manifold shaping makes the coefficient vectors highly clusterable. Second, we run K-means on the generated coefficients, replacing per-edge weights with codebook centroids indexed with compact codes. Finally, we discard the meta-learner and embeddings, and lightly fine-tune the centroids to recover any accuracy loss. Since each centroid stores an entire coefficient vector, the codebook amortizes over many scalars, yielding a much higher compression factor for KANs than for MLPs at the same number of clusters.
@@
We validate MetaCluster on two model families consisting of a fully-connected KAN (Liu et al., 2024b) and a convolutional KAN (ConvKAN) (Bodner et al., 2024; Drokin, 2024). For each of these models, we test the efficacy of using B-Splines, radial basis functions (RBFs), and Gram polynomials as the bases (Liu et al., 2024b; Li, 2024; SS et al., 2024). To further verify the robustness of our approach, we validate it across MNIST (Deng, 2012), CIFAR-10, CIFAR-100 (Krizhevsky, 2009), and equation modeling. From our experiments, we find that MetaCluster achieves up to $80\times$ and $124.1\times$ reductions in parameter storage for image classification and equation modeling, respectively, relative to the uncompressed KAN without degrading accuracy. We also find it is robust across various architectures and datasets, and ablations confirm that enforcing a low-dimensional manifold is key to high-quality clustering.
@@
    We identify the potential of weight-sharing for reducing the KANs memory footprint and, to our knowledge, provide the first effective weight-sharing method tailored to KANs.
@@
    We provide extensive image classification and equation modeling experiments demonstrating up to $80\times$ and $124.1\times$ memory reduction, respectively, with no loss in accuracy, along with extensive ablations.
@@ 2 Preliminaries and Motivation
2.1 Kolmogorov-Arnold Networks
KANs have gained attention as an alternative to conventional MLPs (Liu et al., 2024b). Their design is motivated by the Kolmogorov–Arnold representation theorem, which guarantees that any continuous multivariate function on a bounded domain, $f:[0,1]^{n}\rightarrow\mathbb{R}$, can be expressed as a finite sum of compositions of univariate continuous functions, $\phi_{q,p}:[0,1]\rightarrow\mathbb{R}$ and $\phi_{q}:\mathbb{R}\rightarrow\mathbb{R}$ such that
@@ 2.1 Kolmogorov-Arnold Networks
While Equation 1 captures the classical theoretical form, Liu et al. (2024b) demonstrates a practical generalization that permits arbitrary width and depth tailored to the task. We can express such a formulation with $L$ layers as
$f(\mathbf{x})$ $=(\Phi_{L}\circ\Phi_{L-1}\circ...\circ\Phi_{1})\mathbf{x},$
$\Phi_{l}$ $=\begin{bmatrix}\phi_{l,1,1}(\,\cdot\,)&\cdots&\phi_{l,1,n_{l}}(\,\cdot\,)\\ \vdots&\ddots&\vdots\\ \phi_{l,n_{l+1},1}(\,\cdot\,)&\cdots&\phi_{l,n_{l+1},n_{l}}(\,\cdot\,)\end{bmatrix}$ (2)
In Equation 2.1, we let $n_{l}$ denote the number of inputs to layer $l$, with inputs $x_{l,i}$. Each edge from input $i$ to output $j$ in layer $l$ is equipped with a learnable univariate activation $\phi_{l,j,i}(\cdot)$, for $i\in[1,n_{l}]$ and $j\in[1,n_{l+1}]$.
@@
From its increased representation power compared to the MLP, the KAN has offered impressive results on equation modeling tasks for scientific applications (Li, 2024; Li et al., 2025; Coffman and Chen, 2025; Koenig et al., 2024) and has even started to extend its reach into computer vision (Raffel and Chen, 2025; Yang and Wang, 2024). Despite these gains, widespread adoption in modern large-scale architectures has been hindered by memory inefficiency (Yu et al., 2024). The root cause is structural: each KAN edge carries a vector of basis coefficients (e.g., B-spline weights), whereas an MLP edge carries a single scalar. Ignoring biases, an arbitrary MLP with $L$ layers will possess $\sum_{l=0}^{L-1}(n_{l}\times n_{l+1})$ parameters, whereas a KAN will possess $\sum_{l=0}^{L-1}(n_{l}\times n_{l+1})\times(|\mathbf{w}|)$ parameters. Thus, for identical topologies, a KAN is approximately $|\mathbf{w}|$ times larger than its MLP counterpart.
2.2 KAN Weight Sharing Motivations and Challenges
These observations motivate a compression strategy that targets the dimensionality of the vector of basis coefficients. Weight sharing is one method that reduces dimensionality directly by clustering parameters into a small codebook and storing compact indices. In its classical form for MLPs, one applies K-means to the set of scalar weights $\mathbf{W}={w_{1},w_{2},\dots,w_{n}}$, obtaining centroids $\mathbf{C}={c_{1},c_{2},\dots,c_{k}}$ and assignments that minimize within-cluster sum of squares (Han et al., 2015):
@@ 2.2 KAN Weight Sharing Motivations and Challenges
While weight sharing is promising to save substantial storage for KANs, directly applying it to KANs for model compression is nontrivial. For instance, since each edge on a KAN is defined by $|\mathbf{w}|$ weights rather than $1$ as in the MLP, the dimensionality of the points we must cluster is far greater. In increasing the dimensionality, the vector space becomes sparser and points become more equally spaced apart, making clustering points more difficult (Beyer et al., 1999; Donoho and others, 2000). Therefore, to effectively cluster the KAN activations, we require a fundamentally different method for transforming the high-dimensional vector into a more manageable space.
@@ 3 Methods
This section presents MetaCluster, a three-stage framework that makes per-edge KAN activation coefficients amenable to clustering and compression. First, we train a single meta-learner that maps compact embeddings to full activation-coefficient vectors, thereby forcing them onto a task-aligned low-dimensional manifold. Next, we apply K-means to those vectors and record, for each edge, the centroid to which it belongs in that layer. Finally, we discard the meta-learner and embeddings, replace per-edge weights with their assigned centroids via a lookup table, and briefly fine-tune the network to recover any lost accuracy.
@@ 3.1 Manifold Learning through a Meta-learner
To demonstrate the manifold-shaping effect of the meta-learner, we train a two-layer KAN with a B-spline basis on CIFAR-10 (Krizhevsky, 2009) for one epoch under three settings: (i) a meta-learner with $d_{\text{emb}}=1$, (ii) a meta-learner with $d_{\text{emb}}=2$, and (iii) a baseline KAN without a meta-learner. We collect all per-edge activation coefficient vectors $\mathbf{w}_{i}$} and visualize them with t-SNE. Each point corresponds to $\mathbf{w}_{i}\in\mathbb{R}^{|\mathbf{w}|}$, where $|\mathbf{w}|=G+k+1$ with $G=5$ and $k=3$, yielding $|\mathbf{w}|=9$. As shown in Figure 1(a), when $d_{\text{emb}}=1$ the model organizes coefficients along an effectively one-dimensional manifold. Increasing to $d_{\text{emb}}=2$ produces the well-structured two-dimensional sheet shown in Figure 1(b), indicating that the meta-learner concentrates variability onto task-relevant directions. In contrast, the baseline KAN provided in Figure 1(c) yields a diffuse cloud without coherent low-dimensional structure, which hampers downstream clustering. These visualizations substantiate our premise: shaping coefficient vectors via a low-dimensional embedding makes them markedly more clusterable, a property we later translate into the higher compression at fixed accuracy shown in Section 4.2.
@@
3.2 KAN Activation Compression with MetaClustering
@@ 3.4 Storage Analysis
As in Equation 5, $n$ denotes the number of connections, $k$ the number of centroids, and $b$ the number of bits used to represent each edge. Notably, the term $|\mathbf{w}|$ in the denominator reduces the relative contribution of $n\log_{2}(k)$ to the overall storage. Intuitively, this happens because each centroid stores more weight information ($|\mathbf{w}|$ entries per centroid), so the overhead of indexing and storing the centroids becomes proportionally smaller. As a result, KAN benefits more from the compression scheme than MLP.
@@ 4 Experiments
We provide extensive experiments to showcase the efficacy of MetaCluster. In Section 2, we present the main results of the fully connected and convolutional architectures on MNIST, CIFAR10, and CIFAR100 (Deng, 2012; Krizhevsky, 2009). Next, in Section 4.2, we provide the ablation study for our approach, covering the influence of the KAN meta-learner embedding size, basis coefficient vector sizes, and the cluster count on CIFAR10 (Krizhevsky, 2009). In Section 5 we report the computational cost incurred from MetaCluster. Finally, in Section 4.4, we provide high-dimensional equation modeling results with MetaCluster.
4.1 Image Classification Results
(a) The sample clusters from the first-layer of FastKAN.
(b) The sample clusters from the first-layer of MetaFastKAN.
Figure 2: A set of sample clusters showcasing the edge functions for each cluster (blue) and the centroid functions (red) for the first layer of the fully-connected (a) FastKAN and (b) MetaFastKAN models on Cifar-10.
For all image classification experiments, we did a 90/10 training/validation split of the training data. We apply data augmentation to the train set for the convolutional architecture, following an identical procedure to Drokin (2024). We compare our approach across 24 model schemes. The base model for each scheme consists of KAN (B-spline basis) (Liu et al., 2024b), FastKAN (RBF basis) (Li, 2024), and GramKAN (Gram polynomial basis)(Drokin, 2024). In our naming conventions, we include Cluster to designate a clustered model, Meta to designate a model using a meta-learner, a design that originates from Zhao et al. (2025), and Conv to designate a model with a convolutional architecture (Bodner et al., 2024; Drokin, 2024). The fully-connected and convolutional models were clustered using 16 and 256 clusters, respectively. Each result is reported as an average of 5 runs. We also provide standard error bars. For our image classification experiments, a full description of our hyperparameters is given in Appendices A.1.1 and A.1.2.